OpenStudy (anonymous):

if h(x) = −3x − 10, find h(−3).

OpenStudy (anonymous):

Plug in -3 for X: -3(-3) - 10

OpenStudy (anonymous):

h(-3)=-3(-3)-10=9-10=-1

OpenStudy (anonymous):

Thanks